# Is your stochastic signal really detectable?

### v1

@@ Abstract
Introduction –  Stochastic signals arising from the incoherent superposition of numerous faint sources generated by some physical process are the target for many different types of scientific experiment. One example is the stochastic gravitational wave background (SGWB) sourced by either astrophysical systems or cosmological processes. At nHz frequencies, the first evidence for a SGWB has recently been reported by multiple collaborations [1, 2, 3, 4]. SGWB inference presents a number of challenges, ubiquitous across the spectrum. In particular, it entails disentangling the contribution from the target physical process from that of the instrumental noise, about which knowledge is usually limited, and often requires distinguish the contributions from multiple overlapping components.
@@
Modeling –  Throughout the paper, we assume signal and noise as realizations of two independent, Gaussian, and stationary processes. We denote with $\mathbf{d}=\{\mathbf{d}_{1},\dots,\mathbf{d}_{p}\}$ the datastreams of $p$ detectors in the time domain. These datastreams can represent outputs from a variety of detectors, which for gravitational wave detectors could be Time-Delay Interferometry (TDI) channels from space-based interferometers [12], pulsar timing residuals [13], or individual measurements from ground-based detectors [14]. We divide the total observation time into $n_{c}$ adjacent segments, each of duration $T=T_{\mathrm{obs}}/n_{c}$. We assume the datastreams to be uniformly sampled, with a cadence $\Delta t$, and denote each segment by ${\mathbf{d}}_{(c)}=(\mathbf{d}_{(c){{1}}},\dots,\mathbf{d}_{(c){{p}}})$. We use the segments to construct coarse-grained data, and define the averaged periodograms in the frequency domain as
@@
$\Sigma(f)=\frac{1}{n_{c}}\left(\Sigma_{n}(f)+\Sigma_{\mathrm{GW}}(f)\right),$ (2)
@@
$\!\mathcal{L}(Y(f)|\Sigma(f))\!\propto\!\!\frac{|Y(f)|^{n_{c}-p}}{|\Sigma(f)|^{n_{c}}}\exp\left[-\mathrm{tr}(\Sigma^{-1}(f)Y(f))\right]$ (3)
where tr denotes the trace operator. To evaluate model evidences, priors must be specified and marginalized over as follows
${\cal Z}(Y)=\int{\mathrm{d}}\Sigma{\cal L}(Y\mid\Sigma)\pi(\Sigma),$ (4)
where we drop the frequency dependence of the matrices for brevity. We choose priors conjugate to the likelihood in Eq. 3, such that upon inference the posterior preserves the same functional form. We assume a complex inverse-Wishart prior $\mathcal{CW}^{-1}$ on $\Sigma$, with $\nu$ degrees of freedom and scale matrix $\Psi$.
$\Sigma\mid\Psi,\nu$ $\sim{\cal CW}^{-1}(\Psi,\nu)$ (5)
$\pi(\Sigma|\Psi,\nu)$ $\propto|\Psi|^{\nu}|\Sigma|^{-(\nu+p)}\exp{\left[-\mathrm{tr}(\Psi\Sigma^{-1})\right]}.$ (6)
The chosen prior yields an expectation value for $\Sigma$
$E[\Sigma]=\frac{\Psi}{(\nu-p)}$ (7)
@@
where $\Sigma_{n}^{0}$ denotes a fiducial estimate of the noise covariance matrix, and $\vec{\lambda}$ are the parameters describing the SGWB power spectral density (e.g., the amplitude and slope for a power-law signal). As a consequence, the mean of the prior distribution is the expected $\Sigma$ in Eq. 7.
The value of $\nu$ can be adjusted to represent the scale of our prior uncertainty on $\Sigma$, since the variances $\Sigma_{ii}$ are proportional to $\Psi^{2}_{ii}/\left[(\nu-p)^{2}(\nu-p-1)\right]$. For instance, $\nu\sim 10$ represents almost $10\%$ uncertainty for $p=3$.
As a consequence of conjugacy, the posterior distribution for $\Sigma$ also follows a complex inverse-Wishart distribution with parameters updated according to the observations, which are described by the amount of data observed, $n_{c}$, and the average periodogram, $Y$, from Eq. 1:
$\Sigma\mid Y,\Psi,\nu$ $\sim\mathcal{CW}^{-1}(\Psi+Y,\nu+n_{c}),$ (9)
$p(\Sigma|Y,\Psi,\nu)$ $\!\propto\!\frac{|\Psi|^{\nu}|Y|^{n_{c}-p}}{|\Sigma|^{n_{c}+\nu+p}}\exp\!{\left[-\mathrm{tr}((Y+\Psi)\Sigma^{-1})\right]}$ (10)
@@
where $C$ is a suitable normalization constant, independent on $\Psi,\nu$. Notably, the evidence retains a dependence on the prior hyperparameters $\Psi$ and $\nu$. Reinstating explicitly the dependence on the frequency, the Bayes factor for a given bin reads
@@
using the identity $\ln|M|=\mathrm{tr}(\ln M)$. Using the expansion $\ln\left(\mathbb{I}+Y^{-1}\Psi)=(\sum_{k=1}^{\infty}(-1)^{k+1}(Y^{-1}\Psi)^{k}/k\right)$ at leading order, the log Bayes factor reads
@@
$+\!\frac{(\nu\!+\!n_{c})}{n_{c}-p}\mathrm{tr}(\Sigma^{-1}\Psi_{1}\!-\!\Sigma^{-1}\Psi_{0})\!+\!\mathcal{O}(Y^{-1}\Psi)$ (17)
@@
where we used the relation $E\mathrm{[Y^{-1}]}=\Sigma^{-1}/(n_{c}-p)$ [15]. Even though Is your stochastic signal really detectable? can be expanded by including higher order moments from the complex inverse-Wishart distribution [16], we provide here results at leading-order only. As expected in the case of $\Sigma^{0}_{n}\gg\Sigma_{\rm GW}$, the expectation value of the log-Bayes factor goes to 0. Instead for $\Sigma^{0}_{n}\ll\Sigma_{\rm GW}$, the expectation value is
$\nu\mathrm{tr}\left[\ln(\Sigma_{\mathrm{GW}}(\Sigma_{n}^{0})^{-1})\right]-\frac{(\nu+n_{c})(\nu-p)}{n_{c}-p}\mathrm{tr}\!\left[\mathbb{I}\right]$ (19)
Results –  We show here a concrete example by considering the SGWB detection problem for LISA. We assume three datastreams, i.e., $p=3$, each corresponding to a TDI variable. TDI is a post-processing technique that reduces the laser frequency noise in the LISA measurements. Various combinations of TDI variables can be constructed under different approximations to the orbital motion of the LISA satellites [12].
@@
In this work, we assume equal and constant armlengths, hence use the $A,E,T$ variables as the datastreams [12]. In doing so, the model for $\Sigma$ simplifies to a diagonal matrix. Under these assumptions, the SNR of a signal with parameters $\vec{\lambda}$ reads
@@
where $R_{i}$ is the sky-averaged response function of LISA to SGWB —summed over the two independent polarizations— and $P_{n,i}$ is the power spectral density (PSD) for the noise in the $i$-th TDI channel. Analytical expressions for $R_{i}$ and $P_{n,i}$ are available in the literature [17].
Figure 3: Dependency of BPLS on the observing time. (Left panel) The dashed lines refer to PLSs constructed for a target $\mathrm{SNR}=35$, while solid ones denote the BPLSs at $\log\mathcal{B}=1$ considering $10\%$ prior uncertainties in the covariance matrix. Colors denote three fiducial observation times, $T_{\rm obs}=0.4{\rm yr},2{\rm yr},4{\rm yr}$, as indicated in the legend. (Right panel) Evolution of the amplitude of the PLS (grey) and BPLS (red) as a function of the number of chunks, i.e., the observation time. The amplitudes are relative to the reference frequency of $f_{0}=10^{-2.5}\mathrm{Hz}$. Solid lines represent the best fit of data points to functions $\propto(n_{c})^{-1/2}$ (grey) and $\propto(n_{c}-p)/(n_{c}+\nu)$ (red), as predicted in the main text.
@@
In the analysis presented so far, we assumed a nominal mission duration of four years [10]. To explore the dependence of the BPLS on the observation time, we vary $T_{\mathrm{obs}}$ by adjusting the number of chunks $n_{c}$ at fixed chunk duration, while keeping the prior uncertainty level fixed at $\nu_{10}$. We focus on the dependency on $T_{\rm obs}$ of the BPLS at a single significance $\log\mathcal{B}=1$, which closely matches the PLS with $\mathrm{SNR}=35$ over the nominal mission duration $T_{\mathrm{obs}}=4\mathrm{yr}$. We illustrate our findings in the left panel of Fig. 3. Notably, the PLS shows a stronger dependency on time than the BPLS. At fixed frequency and SNR, the PLS level is inversely proportional to the square root of the observation time, i.e., $\propto 1/\sqrt{T_{\mathrm{obs}}}$. Instead, the BPLS depends non-trivially on $T_{\rm obs}$: from the expansion in Eq. 18, we observe that $E\left[\ln\mathcal{B}\right]\propto(\nu+n_{c})/(n_{c}-p)=(\nu T+T_{\mathrm{obs}})/(T_{\mathrm{obs}}-pT)$. A fit of this form agrees well with the numerical data (see right panel of Fig. 3). The BPLS does in fact show the expected $1/\sqrt{T_{\mathrm{obs}}}$ behaviour, but only in the regime $n_{c}\ll\nu$. This can be derived by considering the $\nu\rightarrow\infty$ limit of Eq. (12). The physical interpretation is that we are able to measure the total noise with a certain precision, which increases as the number of segments increases. For small observation times, the uncertainty in the measured background is large compared to the prior uncertainty and so the latter is irrelevant and the BPLS behaves like the PLS. As the number of segments increases the precision on the total noise improves and there is a transition into a regime in which the prior uncertainty dominates. The precision to which the background amplitude can be determined thus tends to a constant, as seen in the figures.
@@
Conclusion –  Inferring the spectral shape of an SGWB is a non-trivial task. However, in the idealized case of power-law signals, their detectability can be conveniently cast in a simple form. In this work, we addressed some of the shortcomings of the concept of PLS presented in Ref. [5]. In particular, we used Bayesian evidences to define detectability, which allows a marginalization over the uncertain noise properties of the instrument to be included in the definition. We provided a robust method to straightforwardly compute this new BPLS. This curve identifies detectable signal up to a given Bayes factor threshold. To facilitate usage of our results, data behind all figures are available at Ref [19].
@@
We applied our method to the LISA problem under idealized mission conditions. Considering more realistic scenarios, such as correlations between datastreams, introduces additional complexity to the analysis. Nonetheless, such complexity, and possibly parameterized model uncertainties across the spectrum, is expected improve the SGWB detectability. However, the proposed approach remains robust even when a non-diagonal covariance matrix is considered. The BPLS can be extended directly to other signal shape. In that case, we have to construct the $n-1$-dimensional surface at a given Bayes factor, where $n$ denotes the number of parameters describing the template. We then define the Bayesian sensitivity curve by taking the maximum of the PSD over this surface at each frequency.
The proposed methodology is versatile and can be extended to other GW observatories and contexts where stochastic signals are targeted, making it a powerful and flexible tool for the SGWB community.
@@
Acknowledgments –  Data supporting this paper are publicly released at Ref. [19]. The authors thank C. J. Moore, and A. Sesana for useful comments. R.B. is supported by MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. R.B. is supported by Italian Space Agency Grant “Phase A activities for the LISA mission” No. 2017-29-H.0. Computational work was performed at Bicocca’s Akatsuki cluster (B Massive funded).

### v2

@@ Abstract
I Introduction
Stochastic signals arising from the incoherent superposition of numerous faint sources generated by some physical process are the target for many different types of scientific experiment. One example is the stochastic gravitational wave background (SGWB) sourced by either astrophysical systems or cosmological processes. At nHz frequencies, the first evidence for a SGWB has recently been reported by multiple collaborations [1, 2, 3, 4]. SGWB inference presents a number of challenges, ubiquitous across the spectrum. In particular, it entails disentangling the contribution from the target physical process from that of the instrumental noise, about which knowledge is usually limited, and often requires distinguish the contributions from multiple overlapping components.
@@ I Introduction
II Modeling
As previously mentioned, there are features that can be incorporated into the analysis to better characterize a stochastic signal and enhance detection. Throughout this paper, we assume that both signal and noise are realizations of independent, Gaussian, and stationary processes, representing the easiest yet ’worst-case’ scenario for studying the detectability of a stochastic signal. We denote with $\mathbf{d}=\{\mathbf{d}_{1},\dots,\mathbf{d}_{p}\}$ the datastreams of $p$ detectors in the time domain. These datastreams can represent outputs from a variety of detectors, which for gravitational wave detectors could be Time-Delay Interferometry (TDI) channels from space-based interferometers [12], pulsar timing residuals [13], or individual measurements from ground-based detectors [14]. We divide the total observation time into $n_{c}$ adjacent segments, each of duration $T=T_{\mathrm{obs}}/n_{c}$. We assume the datastreams to be uniformly sampled, with a cadence $\Delta t$, and denote each segment by ${\mathbf{d}}_{(c)}=(\mathbf{d}_{(c){{1}}},\dots,\mathbf{d}_{(c){{p}}})$. We use the segments to construct coarse-grained data, and define the averaged periodograms in the frequency domain as
@@ II Modeling
$\Gamma(f)=\frac{1}{n_{c}}\left(\Sigma_{n}(f)+\Sigma_{\mathrm{GW}}(f)\right),$ (2)
@@
$\mathcal{L}(Y(f)|\Gamma(f))\!\propto\!\!\frac{|Y(f)|^{n_{c}-p}}{|\Gamma(f)|^{n_{c}}}\exp\left[-\mathrm{tr}(\Gamma^{-1}(f)Y(f))\right]$ (3)
where tr and $|\cdot|$ denote respectively the trace and the determinant operator. To evaluate model evidences, priors must be specified and marginalized over as follows
${\cal Z}(Y)=\int{\mathrm{d}}\Gamma{\cal L}(Y\mid\Gamma)\pi(\Gamma),$ (4)
where we drop the frequency dependence of the matrices for brevity. We choose priors conjugate to the likelihood in Eq. 3 for two main reasons: first, upon inference the posterior preserves the same functional form of the prior; second, as we shall see below, the free parameters available yield enough flexibility to choose a reference expectation value and variance for $\Gamma$. We assume a complex inverse-Wishart prior $\mathcal{CW}^{-1}$ on $\Gamma$, with $\nu$ degrees of freedom and scale matrix $\Psi$.
$\Gamma\mid\Psi,\nu$ $\sim{\cal CW}^{-1}(\Psi,\nu)$ (5)
$\pi(\Gamma|\Psi,\nu)$ $\propto|\Psi|^{\nu}|\Gamma|^{-(\nu+p)}\exp{\left[-\mathrm{tr}(\Psi\Gamma^{-1})\right]}.$ (6)
The chosen prior yields an expectation value for $\Gamma$
$E[\Gamma]=\frac{\Psi}{(\nu-p)}$ (7)
@@
where $\Sigma_{n}^{0}$ denotes a fiducial estimate of the noise covariance matrix, and $\vec{\lambda}$ are the parameters describing the SGWB power spectral density (e.g., the amplitude and slope for a power-law signal). As a consequence, the mean of the prior distribution is the expected $\Gamma$ in Eq. 7. The value of $\nu$ can be adjusted to represent the scale of our prior uncertainty on $\Gamma$, since the variances $\mathrm{Var(}\Gamma_{ii})$ are proportional to $\Psi^{2}_{ii}/\left[(\nu-p)^{2}(\nu-p-1)\right]$. For instance, $\nu\sim 10$ represents almost $10\%$ uncertainty for $p=3$.
In realistic data-analysis scenarios, the parameter $\vec{\lambda}$ must be inferred, hence a prior on it must be placed. On the contrary following the PLS construction, here we will explore it systematically. The posterior distribution for $\Gamma$ follows a complex inverse-Wishart distribution with parameters updated according to the observations, which are described by the amount of data observed, $n_{c}$, and the average periodogram, $Y$, from Eq. 1:
$\Gamma\mid Y,\Psi,\nu$ $\sim\mathcal{CW}^{-1}(\Psi+Y,\nu+n_{c}),$ (9)
$p(\Gamma|Y,\Psi,\nu)$ $\!\propto\!\frac{|\Psi|^{\nu}|Y|^{n_{c}-p}}{|\Gamma|^{n_{c}+\nu+p}}\exp\!{\left[-\mathrm{tr}((Y+\Psi)\Gamma^{-1})\right]}$ (10)
@@
where $C$ is a suitable normalization constant, independent on $\Psi,\nu$. Notably, the evidence retains a dependence on the prior hyperparameters $\Psi$ and $\nu$. Reinstating explicitly the dependence on the frequency, the Bayes factor for a given bin—between the presence ($H_{1}$) and absence ($H_{0}$) of GW signal hypotheses— can be written as:
@@
using the identity $\ln|M|=\mathrm{tr}(\ln M)$. We now use the expansion $\ln\left(\mathbb{I}+Y^{-1}\Psi)=(\sum_{k=1}^{\infty}(-1)^{k+1}(Y^{-1}\Psi)^{k}/k\right)$, which holds under the assumption that $Y^{-1}\Psi\ll 1$. Since $\Psi\propto(\nu-p)/n_{c}$, this condition corresponds to the limit where $\nu-p$ is small and $n_{c}$ is large. Practically, this represents a scenario characterized by significant uncertainty in the process and a long observation period. At leading order, the log Bayes factor reads
@@
$+\!\frac{(\nu\!+\!n_{c})}{n_{c}-p}\mathrm{tr}(\Gamma^{-1}\Psi_{1}\!-\!\Gamma^{-1}\Psi_{0})\!+\!\mathcal{O}(Y^{-1}\Psi)$ (17)
@@
where we used the relation $E\mathrm{[Y^{-1}]}=\Gamma^{-1}/(n_{c}-p)$ [15]. Even though Section II can be expanded by including higher order moments from the complex inverse-Wishart distribution [16], we provide here results at leading-order only. As expected in the case of $\Sigma^{0}_{n}\gg\Sigma_{\rm GW}$, the expectation value of the log-Bayes factor goes to 0. Instead for $\Sigma^{0}_{n}\ll\Sigma_{\rm GW}$, the expectation value is
$\nu\mathrm{tr}\left[\ln(\Sigma_{\mathrm{GW}}(\Sigma_{n}^{0})^{-1})\right]-\frac{(\nu+n_{c})(\nu-p)}{n_{c}-p}\mathrm{tr}\!\left[\mathbb{I}\right]\,.$ (19)
III Results
We show here a concrete example by considering the SGWB detection problem for LISA. Specifically, we consider a four-year mission duration. To reduce the computational cost we restrict our analysis to the frequency range $\left[10^{-4},10^{-2}\right]\rm Hz$. We assume three datastreams, i.e., $p=3$, each corresponding to a TDI variable. TDI is a post-processing technique that reduces the laser frequency noise in the LISA measurements. Various combinations of TDI variables can be constructed under different approximations to the orbital motion of the LISA satellites [12].
@@ III Results
In this work, we assume equal and constant armlengths, hence use the $A,E,T$ variables as the datastreams [12]. In doing so, the model for $\Gamma$ simplifies to a diagonal matrix. Under these assumptions, the SNR of a signal with parameters $\vec{\lambda}$ reads
@@
where $R_{i}$ is the sky-averaged response function of LISA to SGWB —summed over the two independent polarizations— and $P_{n,i}$ is the power spectral density (PSD) for the noise in the $i$-th TDI channel. Analytical expressions for $R_{i}$ and $P_{n,i}$ are available in the literature [17, 18, 19].
Figure 3: Dependency of BPLS on the observing time. (Left panel) The dashed lines refer to PLSs constructed for a target $\mathrm{SNR}=35$, while solid ones denote the BPLSs at $\log\mathcal{B}=1$ considering $10\%$ prior uncertainties in the covariance matrix. Colors denote three fiducial observation times, $T_{\rm obs}=0.4{\rm yr},2{\rm yr},4{\rm yr}$, as indicated in the legend. (Right panel) Dependency of the PLS (grey) and BPLS (red) value at a reference frequency $f_{0}=10^{-2.5}\mathrm{Hz}$ as a function of the number of chunks, i.e., the observation time. Dashed (solid) curve represent the best fit of 10 PLS (BPLS) values with as expected from Eq. (20) and Eq. (18), respectively.
It is important to note that Eq.20 is valid in the regime where the noise dominates over the signal. Otherwise, more robust statistics need to be considered as in [20].
@@
[…]
IV Conclusion
@@
Inferring the spectral shape of an SGWB is a non-trivial task. However, in the idealized case of power-law signals, their detectability can be conveniently cast in a simple form. In particular, we employed Bayesian evidence as a detectability metric. By leveraging prior conjugacy, we naturally incorporated noise uncertainty and marginalized over them. In analogy with Ref. [5], we developed a robust method to efficiently compute the BPLS. This curve identifies detectable signal up to a given Bayes factor threshold. To facilitate usage of our results, data behind all figures are available at Ref [22].
@@
We applied our method to the LISA problem under idealized mission conditions. Considering more realistic scenarios, such as correlations between datastreams, introduces additional complexity to the analysis. Nonetheless, such complexity, and possibly parameterized model uncertainties across the spectrum, is expected improve the SGWB detectability. However, the proposed approach remains robust even when a non-diagonal covariance matrix is considered. The BPLS can be extended directly to other signal shape. In that case, we have to construct the n1n-1-dimensional surface at a given Bayes factor, where $n$ denotes the number of parameters describing the template. We then define the Bayesian sensitivity curve by taking the maximum of the PSD over this surface at each frequency. The proposed methodology is versatile and can be extended to other GW observatories and contexts where stochastic signals are targeted, making it a powerful and flexible tool for the SGWB community.
Acknowledgements.
@@
Data supporting this paper are publicly released at Ref. [22]. The authors thank C. J. Moore, and A. Sesana for useful comments. R.B. is supported by MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. R.B. is supported by Italian Space Agency Grant “Phase A activities for the LISA mission” No. 2017-29-H.0. Computational work was performed at Bicocca’s Akatsuki cluster (B Massive funded).